OpenStudy (anonymous):

how do you do 11(c-4)+9=2+c-37 step by step because i have tried it over and over and still cant get the right answer

OpenStudy (anonymous):

11c-44+99=c-35 10c=-90 c=-9 actually which part do u stuck on?

OpenStudy (anonymous):

the whole thing

OpenStudy (anonymous):

but do u understand now?

OpenStudy (anonymous):

no where did 44 and 99 and 35 come in at

OpenStudy (anonymous):

11(c-4) when u are solving this bracket, u are actually multiplying 11 into c and -4 so it will be 11c and -4*11 = -44

OpenStudy (anonymous):

well i just realized that i type 99 instead of 9 - -wait

OpenStudy (anonymous):

11c-44+9=c-35 10c=-70 c=-7 this should be the correct one...

OpenStudy (anonymous):

and u got -35 because u are adding 2 into -37, that's from your own equation, the right hand side of it.

OpenStudy (anonymous):

ok where did 44 come from

OpenStudy (anonymous):

it said -7 is wrong answer

OpenStudy (anonymous):

you distribute the 11 into the C and -4 and get 11c-44+9 on the left, which simplifies to 11c-35. You add 2 and -37 together on the right and get C -35. So you have 11c-35 on the left and c-35 on the right. So add 35 on both sides and you get 11c=c. and the only way that 11 times SOMETHING = SOMETHING is if that something is 0.

OpenStudy (anonymous):

omg- - i don't even know what i am typing- - the answer should be 0 orz

OpenStudy (anonymous):

cuz at last u will get 10c = 0 which means c=0.. i just messed it up, sorry:S

OpenStudy (anonymous):

its ok and yes 0 is the answer thanks